# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v3

@@ 4 Experimental Setup
For our experiments, we evaluate our approach across three programming languages—Python, C++, and Java—using four benchmark datasets: HumanEval+ and MBPP+ (Liu et al., 2023b) for Python, and HumanEvalPack (Muennighoff et al., 2024) for C++ and Java. Table 1 presents token-length statistics of the evaluation datasets, computed with the Qwen2.5-Coder-7B tokenizer. The datasets span from short snippets to long sequences, enabling assessment of our method across varying complexity.
@@
We consider the following three watermarking approaches for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP (Guan et al., 2024) from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods. As part of our comparative analysis, we reproduce CodeIP and report its experimental results in Appendix B.
@@
We use Qwen2.5-Coder-7B (Hui et al., 2024) for our experiments. We also report additional results with Llama-3.1-8B in Appendix C, which further demonstrate the generalizability of our approach across different model architectures.
@@
We evaluate watermarking methods across three dimensions: Correctness, Detectability, and Imperceptibility. Correctness uses pass@k (k=1,5), detectability is measured by AUROC with a $z$-score test, and imperceptibility by perplexity computed with StarCoder2-7B. We use the STEM framework to integrate these three axes into a single composite score. For main results (Tables 2 and 3), we adopt the equal-weight configuration $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ as a neutral. We further analyze weight settings with alternative configurations in Appendix D.
@@
During code generation, we apply top-$k$ sampling with $k=50$, restricting the candidate token pool to the top 50 tokens with the highest logits and normalizing their probabilities before sampling. The temperature, which controls the sharpness of the token distribution, is fixed at 1.0. We set $\gamma=0.5$ and $\delta=1.0$ for the MBPP+ dataset, while for the HumanEval+ and HumanEvalPack (C++, Java) datasets, we set $\gamma=0.5$ and $\delta=0.5$. We conduct all experiments on an NVIDIA A6000 GPU.
@@ 5 Experiments and Results
We formulate three key research questions and conduct a thorough analysis to addressing them:
@@
These research questions frame our analysis of STONE along three key aspects: 1) Functionality Preservation: verifying that watermark insertion does not alter program behavior; 2) Balanced Performance under STEM: evaluating whether STONE achieves a well-balanced trade-off among correctness, detectability, and imperceptibility by the STEM metric; and 3) Performance Overhead: measuring the computational cost of watermark insertion and detection.
@@
We examine whether STONE embeds watermarks without affecting the functional correctness of generated code. Tables 2 and 3 show that STONE improves correctness by 7.57% over SWEET across all benchmarks. The token-level analysis in Appendix E indicates that the entropy-based token selection method used by SWEET often chooses syntax-related elements—such as delimiters, whitespace, keywords, types, and operators—that are essential for correct execution. For example, modifying delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can cause parsing errors, and changing operators or keywords (e.g., replacing ‘+’ with ‘-’ or ‘True’ with ‘False’) can alter program logic or control flow. Even under the optimal configuration, syntax tokens account for 12.6% of the tokens selected for watermarking in the HumanEval+ dataset, with delimiters and whitespace forming the majority. This inclusion of syntax elements explains the observed reduction in correctness for SWEET. In contrast, STONE excludes syntax tokens and embeds watermarks only into non-syntactic elements, which prevents structural and logical errors during execution and allows the generated code to preserve its functional behavior.
@@
We examine whether STONE achieves balanced performance across correctness, detectability, and imperceptibility. Baselines such as KGW and EWD watermark every token, and this process dilutes the watermark signal and lowers AUROC scores, while SWEET modifies high-entropy tokens including syntax elements, improving detectability but harming fluency. In contrast, STONE embeds watermarks only in non-syntax tokens, and this design preserves syntactic integrity while keeping the signal distinct. This selective design enables STONE to achieve higher detection accuracy without compromising the imperceptibility of code.
Imperceptibility results reinforce this finding. KGW and EWD achieve reasonable fluency by uniformly adjusting token probabilities, but their adjustments can still introduce subtle artifacts. SWEET often produces more noticeable distributional shifts because it alters syntax tokens. STONE, on the other hand, maintains low perplexity differences by leaving syntax untouched and altering only semantically flexible positions. As a result, STONE sustains imperceptibility while preserving correctness and robustness in detection. This balanced performance is captured by the STEM score under the equal-weight configuration, which provides a neutral evaluation across correctness, detectability, and imperceptibility. Unlike other watermarking methods that exhibit clear trade-offs among these criteria, STONE achieves consistently superior overall performance. A comprehensive grid analysis across 66 weight settings (Appendix D) demonstrates that this advantage remains stable and does not depend on any specific weighting setting.
@@
Table 4: Comparison of total insertion and detection time (measured in seconds) across different watermarking methods.
@@
We analyze the computational overhead of watermarking by comparing insertion and detection time. Insertion time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark. Table 4 presents a summary of the results for each method. While all approaches exhibit comparable insertion times, STONE achieves detection speeds that are, on average, 86% faster than those of SWEET and EWD. This improvement stems from differences in the detection process: both SWEET and EWD rely on entropy-based method, which involves computing token probability distributions and leads to greater computational overhead. In contrast, STONE eliminates this burden by verifying watermarks using a predefined green list of non-syntax tokens. This design reduces detection time without sacrificing detectability.
@@ 6.1 Trade-off Between Code Quality and Detection Performance
We analyze the trade-off between code quality and watermark detectability as influenced by two watermarking parameters. The parameter $\gamma$ controls the proportion of the vocabulary designated as the green list, while $\delta$ determines the extent to which the selection probability of green tokens is amplified. We conduct an analysis on the HumanEval+ dataset by varying the values of $\delta$ and $\gamma$. Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other. The circular markers represent the pass@1 and detection performance (AUROC) of STONE, while the triangular markers indicate those of SWEET. Among the four regions divided by two black dashed lines, the upper-right quadrant corresponds to the area where code quality and detection performance are both high, indicating a desirable balance between utility and detectability. We observe that while only a single configuration of SWEET falls into this balanced region, multiple configurations of STONE—spanning various combinations of $\gamma$ and $\delta$—lie within it. This suggests that STONE achieves a more balanced trade-off between functionality and watermark detectability compared to SWEET.
@@ 6.2 Robustness under Adversarial Attacks
Although a watermark is embedded in the code, a malicious user may remove it by modifying the code. Therefore, the watermark must remain robust against attacks like code modifications. We assess the resilience of watermarks of STONE against code alterations by applying two types of attacks: (1) a code refactoring attack using a commercial service [^1] https://codepal.ai/code-refactor and (2) a code paraphrasing attack using ChatGPT (GPT-4o) [^2] Prompt: Please paraphrase the following code.. We conduct this analysis on the HumanEval+ dataset. Figure 4 shows the classification performance (i.e., detectability) in distinguishing between watermarked LLM-generated code and human-written code after these attacks are applied to code watermarked by STONE and SWEET. The experimental results show that STONE is more robust than SWEET against both types of attacks. Refactoring reorganizes code while keeping syntax boundaries, so STONE preserves many watermarks inserted into non-syntax tokens. Paraphrasing changes variable names and expressions, which directly correspond to the embedding targets of STONE, and thus decreases detectability. SWEET shows the opposite tendency because refactoring alters its watermarking targets. STONE achieves higher robustness under both attacks because it embeds watermarks into a syntax-filtered token space that remains stable across structural and lexical changes.
Figure 4: We evaluate the robustness of watermarking methods by applying two types of adversarial perturbations—code refactoring and paraphrasing—to watermarked code. Following these attacks, we assess the effectiveness of watermark detection. We compare the performance of watermarks embedded using SWEET and STONE to determine their resilience under both attack scenarios.
@@ 7 Related Work
Post-hoc detection methods (Xu and Sheng, 2024; Shi et al., 2024) identify machine-generated code after generation using statistical cues such as perplexity or syntax differences. In contrast, watermarking embeds provenance signals during generation to enable direct identification. This work focuses on the latter approach, developing a syntax-aware watermarking method for code generation.
Among watermarking approaches, Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness. Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks. We address the limitations observed in SWEET and CodeIP by explicitly skipping watermark insertion for syntax-critical tokens. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead. ACW (Li et al., 2024a) and SrcMarker (Yang et al., 2024) employed watermarking techniques based on code transformations, such as semantic-preserving edits and variable renaming. While these approaches preserved program functionality, the resulting modifications often introduce unnatural code patterns, which can increase the risk of watermark removal by adversaries.
@@ 8 Conclusion
We present STONE, a syntax-aware watermarking method that preserves functional correctness by excluding syntax-critical tokens during watermark insertion. We also propose STEM, a unified evaluation framework integrating correctness, detectability, and imperceptibility for balanced assessment. Experiments on Python, C++, and Java show that STONE outperforms baseline methods without degrading code quality and reduces watermark detection time compared with entropy-based approaches. Evaluation under the STEM framework demonstrates balanced performance across weight configurations, and further analysis confirms that STONE maintains stable and robust performance under code modifications.
@@
While STONE establishes a robust methodology for functionality-preserving watermarking, it has several limitations.
@@
Second, although our evaluation demonstrates resilience to general attacks like paraphrasing and refactoring, a more sophisticated threat model considers an adversary with specific knowledge of the STONE algorithm. Such an adversary may directly attack non-syntactic tokens by systematically renaming all variables and removing comments. This class of attack poses a more significant threat to watermark integrity than general modifications. Developing countermeasures against such targeted, algorithm-aware attacks remains a critical avenue for future research.
Appendix A Syntax Elements in Python, C++, and Java
Table 7 categorize the syntactic elements into five groups: keywords, whitespace, types, delimiters, and operators. The selection of syntactic elements reflects the distinct characteristics of each programming language.
Appendix B Performance Analysis of CodeIP
@@
Table 5: Evaluation of CodeIP on four datasets. Correctness: pass@1; Detectability: AUROC; Imperceptibility: PPL gap (negative means degradation).
@@
CodeIP maintains high detectability, achieving AUROC values above 0.94 across all datasets. This result indicates that grammar-based watermark insertion enables reliable message extraction. However, the strong detection capability is accompanied by severe degradation in correctness and imperceptibility. The pass@1 scores decrease sharply, approaching zero on HumanEval+ and HumanEvalPack-C++, which shows that syntactic validity enforced by the type predictor does not ensure functional accuracy. The large perplexity gap of up to +4.9 indicates that the interaction between watermarking and type-prediction constraints distorts the token distribution, causing deviations from the model’s natural generation pattern.
@@
Table 6 presents the STEM scores of CodeIP under different weight settings. Across all configurations, detectability remains consistently high, while correctness- or imperceptibility-focused STEM scores are significantly lower than those of other methods. Even under weight settings emphasizing correctness or imperceptibility, CodeIP does not achieve balanced performance, showing limited robustness under varying evaluation preferences. These findings demonstrate that maximizing watermark detectability without syntax awareness severely degrades code quality. In contrast, STONE embeds watermarks only within non-syntax tokens and achieves stable performance across correctness, detectability, and imperceptibility as captured by the STEM metric.
@@
Appendix C Experimental Results on Llama-3.1-8B
Tables 8 and 9 summarize the experimental results obtained with the Llama-3.1-8B model on MBPP+ and HumanEval+. STONE achieves the highest STEM scores across all datasets and weight settings, demonstrating that the non-syntax watermarking strategy generalizes well to larger models. EWD shows slightly higher detectability on HumanEval+, which appears to result from an interaction between its uniform insertion strategy and the sharper probability distributions of Llama-3.1. In this setting, the cumulative bias introduced by EWD is more effectively reinforced across long and diverse sequences, leading to a small gain in detectability. Such an effect is not observed under Qwen2.5-Coder-7B, whose flatter distributions yield comparable results between EWD and SWEET. Overall, STONE preserves stronger correctness and achieves a more balanced trade-off, resulting in the highest overall STEM scores.
Table 9: STEM scores on MBPP+ and HumanEval+ (Llama-3.1-8B) under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility. Bold indicates the highest score for each weight configuration.

### v4

@@ 4 Experimental Setup
For our experiments, we evaluate our approach across three programming languages: Python, C++, and Java. We use four benchmark datasets: HumanEval+ and MBPP+ (Liu et al., 2023) for Python, and HumanEvalPack (Muennighoff et al., 2024) for C++ and Java. Table 1 reports token-length statistics of the evaluation datasets, based on the Qwen2.5-Coder-7B tokenizer. The datasets span from short snippets to long sequences, allowing evaluation under varying code lengths and structural complexity.
@@
We consider the following three watermarking approaches for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP (Guan et al., 2024) from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods. As part of our comparative analysis, we reproduce CodeIP and report its experimental results in Appendix A.
@@
We use Qwen2.5-Coder-7B (Hui et al., 2024) for our experiments. We also report additional results with Llama-3.1-8B in Appendix B, which further demonstrate the generalizability of our approach across different model architectures.
@@
We evaluate watermarking methods across three dimensions: Correctness, Detectability, and Imperceptibility. Correctness uses pass@k (k=1,5), detectability is measured by AUROC with a $z$-score test, and imperceptibility by perplexity computed with StarCoder2-7B. We use the STEM metric to integrate these three axes into a single composite score. For the main results (Tables 2 and 3), we adopt an equal-weight configuration $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$. We further analyze weight settings with alternative configurations in Appendix D.
@@
During code generation, we apply top-$k$ sampling with $k=50$, restricting the candidate token pool to the top 50 tokens with the highest logits and normalizing their probabilities before sampling. The temperature, which controls the sharpness of the token distribution, is fixed at 1.0. We set $\gamma=0.5$ and $\delta=1.0$ for the MBPP+ dataset, while for the HumanEval+ and HumanEvalPack datasets (C++ and Java), we set $\gamma=0.5$ and $\delta=0.5$. We conduct all experiments on an NVIDIA A6000 GPU.
@@ 5 Experiments and Results
We pose three research questions and provide a thorough analysis of each:
@@
We address these questions through three analytical dimensions: (1) Functionality Preservation, which verifies that watermark insertion does not alter program behavior; (2) Balanced Performance under STEM, which examines whether STONE maintains a well-balanced trade-off among correctness, detectability, and imperceptibility; and (3) Efficiency, which measures the computational overhead of watermark insertion and detection.
@@
We examine whether STONE embeds watermarks without affecting the functional correctness of generated code. Tables 2 and 3 show that STONE improves correctness by 7.57% over SWEET on average across all benchmarks. The token-level analysis in Appendix E indicates that the entropy-based token selection method used by SWEET often chooses syntax-related elements—such as delimiters, whitespace, keywords, types, and operators—that are essential for correct execution. For example, modifying delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can cause parsing errors, and changing operators or keywords (e.g., replacing ‘+’ with ‘-’ or ‘True’ with ‘False’) can alter program logic or control flow. Even under the optimal configuration, syntax tokens account for 12.6% of the tokens selected for watermarking in the HumanEval+ dataset, with delimiters and whitespace forming the majority. This inclusion of syntax elements explains the observed reduction in correctness for SWEET. In contrast, STONE excludes syntax tokens and embeds watermarks only into non-syntactic elements, which prevents structural and logical errors during execution and allows the generated code to preserve its functional behavior.
@@
We examine whether STONE achieves balanced performance across correctness, detectability, and imperceptibility. Baselines such as KGW and EWD watermark every token, which dilutes the watermark signal and lowers AUROC scores. SWEET modifies high-entropy tokens, including syntax elements, improving detectability but reducing syntactic stability. In contrast, STONE embeds watermarks only in non-syntax tokens, preserving syntactic integrity while keeping the signal distinct. This selective design enables STONE to achieve higher detection accuracy without compromising code imperceptibility.
Imperceptibility results reinforce this finding. KGW and EWD maintain reasonable imperceptibility by uniformly adjusting token probabilities, but their adjustments can still introduce subtle artifacts. SWEET often produces more noticeable distributional shifts because it alters syntax tokens. STONE, in contrast, maintains low perplexity differences by leaving syntax untouched and modifying only non-syntax tokens. As a result, STONE maintains imperceptibility without sacrificing correctness or detection robustness. We quantify this balanced behavior using the STEM metric, which under the equal-weight configuration provides a neutral evaluation across correctness, detectability, and imperceptibility. Unlike other watermarking methods that exhibit clear trade-offs among these criteria, STONE achieves consistently strong overall performance. A comprehensive grid analysis across 66 weight settings (Appendix D) further confirms that this advantage remains stable and does not depend on any specific weighting scheme.
@@
Table 4: Comparison of total insertion and detection time (measured in seconds) across different watermarking methods. The reported measurement is performed at the dataset level.
@@
We analyze the computational overhead of watermarking in terms of insertion and detection time. Insertion time denotes the duration required to generate watermarked code, whereas detection time measures the time needed to verify a watermark. Table 4 summarizes the results for each method. All approaches exhibit comparable insertion times, but STONE achieves detection speeds that are on average 86% faster than those of SWEET and EWD at the dataset level. This improvement results from differences in detection mechanisms. SWEET and EWD employ entropy-based approaches that compute token probability distributions, which substantially increase computational overhead. In contrast, STONE verifies watermarks using a predefined green list of non-syntax tokens, reducing detection time while maintaining reliable detectability.
@@ 6.1 Trade-off Between Code Quality and Detection Performance
We analyze the trade-off between code quality and watermark detectability as influenced by two watermarking parameters. The parameter $\gamma$ controls the proportion of the vocabulary designated as the green list, while $\delta$ determines the extent to which the selection probability of green tokens is amplified. We conduct an analysis on the HumanEval+ dataset by varying the values of $\delta$ and $\gamma$. Figure 3 shows the results in a bubble chart, revealing a clear trade-off between code quality and detectability. The circular markers represent the pass@1 and detection performance (AUROC) of STONE, while the triangular markers indicate those of SWEET. Among the four regions divided by the two black dashed lines, the upper-right quadrant corresponds to configurations where both code quality and detection performance are high, indicating a desirable balance between utility and detectability. We observe that, while only a single configuration of SWEET falls into this balanced region, multiple configurations of STONE—spanning various combinations of $\gamma$ and $\delta$—lie within it. This indicates that STONE achieves a more balanced trade-off between functionality and watermark detectability compared to SWEET.
@@ 6.2 Robustness under Adversarial Attacks
Although a watermark is embedded in the code, a malicious user may remove it by modifying the code. Therefore, the watermark must remain robust against attacks like code modifications. We assess the resilience of watermarks of STONE against code alterations by applying two types of attacks: (1) a code refactoring attack using a commercial service [^1] https://codepal.ai/code-refactor and (2) a code paraphrasing attack using ChatGPT (GPT-4o) [^2] Prompt: Please paraphrase the following code.. We conduct this analysis on the HumanEval+ dataset. Figure 4 shows the classification performance (i.e., detectability) in distinguishing between watermarked LLM-generated code and human-written code after these attacks are applied to code watermarked by STONE and SWEET. The experimental results show that STONE is more robust than SWEET against both types of attacks. Refactoring reorganizes code while keeping syntax boundaries, so STONE preserves many watermarks inserted into non-syntax tokens. Paraphrasing changes variable names and expressions, which directly correspond to the embedding targets of STONE, and thus decreases detectability. SWEET shows the opposite tendency because refactoring alters its watermarking targets. STONE maintains higher robustness under both attacks by embedding watermarks into a syntax-filtered token space that remains stable under structural and lexical modifications. This result confirms that STONE maintains consistent detection robustness across different attack types.
Figure 4: We evaluate robustness by applying code refactoring and paraphrasing attacks to watermarked code and measuring detection performance. Results compare the resilience of SWEET and STONE under both attacks.
@@ 7 Related Work
Post-hoc detection methods (Xu and Sheng, 2024; Shi et al., 2024) identify machine-generated code after generation using statistical cues such as perplexity or syntax differences. In contrast, watermarking embeds provenance signals during generation to enable direct identification (Kirchenbauer et al., 2023; Guan et al., 2024; Park et al., 2025). Our work focuses on the latter approach, developing a syntax-aware watermarking method for code generation.
Among watermarking approaches, Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness. Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks. In contrast to these approaches, our method explicitly excludes syntax-critical tokens from watermark insertion. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead. ACW (Li et al., 2024) and SrcMarker (Yang et al., 2024) employed watermarking techniques based on code transformations, such as semantic-preserving edits and variable renaming. While these approaches preserved program functionality, the resulting modifications often introduce unnatural code patterns, which can increase the risk of watermark removal by adversaries.
@@ 8 Conclusion
We present STONE, a syntax-aware watermarking method that preserves functional correctness by excluding syntax-critical tokens during watermark insertion. We also propose STEM, a unified evaluation metric integrating correctness, detectability, and imperceptibility for balanced assessment. Experiments on Python, C++, and Java show that STONE outperforms baseline methods without degrading code quality and reduces watermark detection time compared with entropy-based approaches. Evaluation under the STEM metric demonstrates balanced performance across weight configurations, and further analysis confirms that STONE maintains stable and robust performance under code modifications.
@@
While STONE provides a robust approach to functionality-preserving watermarking, it has several limitations.
@@
Second, although our evaluation demonstrates resilience to general attacks like paraphrasing and refactoring, a more sophisticated threat model considers an adversary with specific knowledge of the STONE algorithm. Such an adversary may directly attack non-syntactic tokens by systematically renaming all variables and removing comments. This class of attack poses a more significant threat to watermark integrity than general modifications. Developing countermeasures against such targeted, algorithm-aware attacks remains an important direction for future research.
Acknowledgments
This research was supported by the NRF grant (RS-2025-00562134) and the AI Graduate School Program (RS-2020-II201361) funded by the Korean government.
Appendix A Performance Analysis of CodeIP
@@
Table 5: Evaluation of CodeIP on four datasets. Correctness: pass@1; Detectability: AUROC; Imperceptibility: PPL change from non-watermarked to watermarked code.
@@
CodeIP maintains high detectability, achieving AUROC values above 0.94 across all datasets. This result indicates that grammar-based watermark insertion enables reliable message extraction. However, this detection capability is accompanied by substantial drops in correctness and imperceptibility. The pass@1 scores decrease sharply, approaching zero on HumanEval+ and HumanEvalPack-C++, which shows that syntactic validity enforced by the type predictor does not necessarily imply functional correctness. The large perplexity gap of up to +4.9 indicates that the interaction between watermarking and type-prediction constraints can shift the token distribution, leading to deviations from the model’s natural generation pattern.
@@
Table 6 presents the STEM scores of CodeIP under different weight settings. Across all configurations, detectability remains consistently high, while correctness- or imperceptibility-focused STEM scores are significantly lower than those of other methods. Even under weight settings emphasizing correctness or imperceptibility, CodeIP does not achieve balanced performance, showing limited robustness under varying evaluation preferences. These findings demonstrate that maximizing watermark detectability alone severely degrades code quality. In contrast, STONE embeds watermarks only within non-syntax tokens and achieves stable performance across correctness, detectability, and imperceptibility as captured by the STEM metric.
Appendix B Experimental Results on Llama-3.1-8B
Tables 7 and 8 summarize the experimental results obtained with the Llama-3.1-8B model on MBPP+ and HumanEval+. STONE achieves the highest STEM scores across all datasets and weight settings, demonstrating that the non-syntax watermarking strategy generalizes well to larger models. EWD shows slightly higher detectability on HumanEval+, which may be related to differences in the output distribution of Llama-3.1-8B under uniform insertion. In this setting, the cumulative bias introduced by EWD is more effectively reinforced across long and diverse sequences, leading to a small gain in detectability. Such an effect is not observed under Qwen2.5-Coder-7B, whose flatter distributions yield comparable results between EWD and SWEET. Overall, STONE preserves stronger correctness and achieves a more balanced trade-off, resulting in the highest overall STEM scores.
Table 8: STEM scores on MBPP+ and HumanEval+ (Llama-3.1-8B) under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility.
@@
Appendix C Syntax Elements in Python, C++, and Java
Table 9 categorizes the syntactic elements into five groups: keywords, whitespace, types, delimiters, and operators. The selection of syntactic elements reflects the distinct characteristics of each programming language.